# Block-Recurrent Dynamics in ViTs

### v1

@@ Abstract
As Vision Transformers (ViTs) become standard backbones across vision, a mechanistic account of their computational phenomenology is now essential. Despite architectural cues that hint at dynamical structure, there is no settled framework that interprets Transformer depth as a well-characterized flow. In this work, we introduce the Block-Recurrent Hypothesis (BRH), arguing that trained ViTs admit a block-recurrent depth structure such that the computation of the original $L$ blocks can be accurately rewritten using only $k\ll L$ distinct blocks applied recurrently. Across diverse ViTs, between-layer representational similarity matrices suggest few contiguous phases. Yet, representational similarity does not necessarily imply functional similarity. To determine whether these phases reflect genuinely reusable computation, we operationalize our hypothesis in the form of block recurrent surrogates of pretrained ViTs, which we call Recurrent Approximations to Phase-structured TransfORmers (Raptor). Using small-scale ViTs, we demonstrate that phase-structure metrics correlate with our ability to accurately fit Raptor and identify the role of training and stochastic depth in promoting the recurrent block structure. We then provide an empirical existence proof for BRH in foundation models by showing that we can train a Raptor model to recover $96\%$ of DINOv2 ImageNet-1k linear probe accuracy in only 2 blocks while maintaining equivalent computational cost. To provide a mechanistic account of these observations, we leverage our hypothesis to develop a program of Dynamical Interpretability. We find (i) directional convergence into class-dependent angular basins with self-correcting trajectories under small perturbations, (ii) token-specific dynamics, where cls executes sharp late reorientations while patch tokens exhibit strong late-stage coherence reminiscent of a mean-field effect and converge rapidly toward their mean direction, and (iii) a collapse of the update to low rank in late depth, consistent with convergence to low-dimensional attractors.
@@ 1 Introduction
In the last decade, Transformers have become the default neural network architecture across machine learning communities, scaling favorably with data and compute (vaswani2017attention; kaplan2020scalinglawsneurallanguage). In particular, Vision Transformers (ViTs) (dosovitskiy2020image) have become the core architecture used in visual foundation modeling frameworks such as DINOv2 (oquab2023dinov2; darcet2023vision) and CLIP (radford2021learningtransferablevisualmodels); and have come to dominate a wide range of visual tasks, from general visual feature extraction (he2021maskedautoencodersscalablevision; chiu2024fine; yun2025vision), to diffusion (peebles2023scalablediffusionmodelstransformers), image segmentation (kirillov2023segment; liu2024grounding), and video processing (arnab2021vivitvideovisiontransformer; baldassarre2025back). This increasing breadth of use motivates a move from empirical optimization to principled understanding.
@@
Two pressures make this understanding urgent. First, safety-critical deployments (wang2022artificial; alecu2022can) demand mechanisms whose internal computation is inspectable (losch2021semantic), diagnosable (adebayo2020debugging), and verifiable (tjeng2017verifying) rather than opaque. As these models proliferate across domains, the ability to explain (doshivelez2017rigorous; gilpin2018explaining; kim2018interpretability), manipulate, and verify their behavior becomes increasingly essential. Second, from a scientific inference perspective (cichy2019deep), understanding what makes these models work is essential for explaining their success. Their strong performance suggests they have discovered effective strategies, and identifying these strategies could advance our broader understanding of visual intelligence. Independently of any comparison to human vision, the goal is to uncover the principles that make these systems so effective.
@@
One promising path toward such understanding is to search for underlying simplicity. Multiple approaches explore different facets of this simplicity, whether in functional expressivity (montufar2014number; telgarsky2015representation; serra2018bounding; balestriero2018spline), symmetry (cohen2014learningirreduciblerepresentationscommutative; olah2020naturally), or computation (wilson2025deep; goldblum2023no; SCHMIDHUBER1997857; mingard2025deep). Discovering such simplicity principle should improve both development (bronstein2017geometric; frankle2018lottery) and interpretability (bereska2024mechanistic; survey2019metrics; fel2024sparks; ghorbani2017interpretation; fel2023holistic; smilkov2017smoothgrad; sundararajan2017axiomatic; zeiler2013visualizing; templeton2024scaling; bricken2023monosemanticity). Depth offers a concrete place to look for this simplicity. Residual connections have long suggested a link to dynamical systems (liao2020bridginggapsresiduallearning; veit2016residual; greff2016highway; boulch2017sharesnet; Haber_2017), hinting at implicit recurrence even when layers have distinct parameters. This convergent evidence makes plausible a form of algorithmic parsimony (ma2022principles) in which a small set of blocks may be reused across many layers, trading parameters for iterations. Related perspectives support this view (dingle2020generic). More concretely, residual updates invite a discrete-time interpretation of depth (chen2018neural; chalvidal2020go; sander2022residual), attention induces coupled token dynamics (lu2019understanding; geshkovski2023emergence), and in language models contiguous block recurrence has been observed and exploited (geiping2025scalingtesttimecomputelatent; fernando2025transformerdynamicsneuroscientificapproach; dehghani2018universal; tan2023sparse). This renewed interest in recurrent computation (jolicoeur2025less; venkatraman2025recursive) and algorithmic complexity (shaw2025bridging; dingle2020generic) as frameworks for understanding neural network simplicity bias further motivates our investigation.
@@
However, no existing framework characterizes depth in ViTs as representational flow or determines whether apparent phases correspond to functional reuse. Furthermore, vision explainability research (bach2015pixel; fong2017meaningful; novello2022making; muzellec2023gradient; petsiuk2018rise; hedstrom2022quantus; fel2025archetypal; gorton2024missing; kowal2024visual; bau2017network; vilas2023analyzing) has not leveraged dynamical systems analysis to model emergent network structure.
@@ 2 Emergent Phase Structure & the Block-Recurrent Hypothesis
Our investigation starts with a simple experiment: we construct layer-layer similarity matrices by computing the cosine similarity of each token at layer $l$ with the same token at layer $m$. As shown in Figure 1, despite significant differences in tasks, training objectives, and scale, all models exhibit consistent block-wise organization where contiguous layers exhibit high mutual similarity within blocks and lower similarity across block boundaries. This finding is not new and echoes early (and more recent) observations in residual networks (kornblith2019similarityneuralnetworkrepresentations; hoang2025comparative), but raises a fundamental question: does representational similarity reflect deeper computational structure? In fact, representational similarity alone provides no guarantee of functional equivalence. Layers might produce similar representations through entirely different computational pathways, or conversely, functionally equivalent computations might yield representations that appear dissimilar due to linear transformations or noise. The critical question is whether these apparent phases correspond to genuine functional recurrence – that is, whether the same computational operations are being reused across different layers within each phase. We formalize this possibility through the Block-Recurrent Hypothesis:
@@
However, recurrent architectures are notoriously difficult to train: rollouts can drift as small state errors compound over steps, and learning can be unstable because gradients propagated through many recurrent applications often vanish or explode (pascanu2013difficulty; linsley2020stable; de2024griffin). Standard backpropagation through time becomes unstable as recurrence depth increases, and the model must learn to simultaneously handle both the forward dynamics and its own prediction errors in closed loop (williams1989learning). To circumvent these challenges, we will leverage the intermediate layer activations as training targets, enabling a staged approach that combines the stability of teacher forcing with the self-consistency required for autoregressive deployment. More precisely, we introduce a procedure to distill existing Vision Transformers into Recurrent Approximations to Phase-structured TransfORmers (Raptors), using $k$ parameter-tied blocks with repetition counts determined by a max-cut phase discovery algorithm. This approach transforms the abstract hypothesis into a concrete architectural and training framework that can be empirically validated.
This constructive approach requires that Raptor models reproduce the internal activations of the full ViT they approximate, similar to dasgupta2025improving; sanh2019distilbert; shleifer2020pre, not merely mimic the final output [^1] Unlike classical distillation, which typically supervises logits (and occasionally a few intermediate “hints”), we enforce one-to-one alignment of all layers representations across the entire depth for the same inputs. The recurrent surrogate must generate the teacher’s intermediate activations, not just its predictions.. The BRH implies that such reproduction should be possible within tolerance $\varepsilon$, making activation matching a natural training objective. Formally, let $\bm{f}$ be a reference ViT with intermediate activations $\bm{a}_{\ell}(\bm{x})\equiv\bm{f}_{\ell}(\bm{x})\in\mathbb{R}^{t\times d}$ for $\ell=0,\dots,L$, where layer $\ell=0$ denotes the patch encoder and $1\leq\ell\leq L$ refer to transformer layers. Here, $t$ is the number of tokens and $d$ the feature dimension. Let $\bm{B}_{j}$ denote the $j$-th parameter-tied block in our recurrent decomposition. The Raptor approximation produces activations:
@@
For a given number of blocks $k$, we must introduce a practical method to determine the number of recurrent iterations of each block ($n_{k}$); in other words, where the recurrent “phases” of computation begin and end. We accomplish this by casting this “block discovery” process as a weighted max-cut problem goemans1995improved, solved via dynamic programming (see Subsection˜A.1 for details). Specifically, the algorithm seeks to partition depth into contiguous segments by maximizing within-block similarity and minimizing cross-block similarity. We visualize the results of this procedure applied to ViT-B in Figure˜3, demonstrating that the discovered blocks align reasonably with qualitative assessment.
@@
Having operationalized the BRH and demonstrated a method for block discovery, we now turn to the mechansitic origins of this phenomenon: under what conditions does this block-recurrent structure emerge in trained Vision Transformers? To investigate this systematically, we examine small-scale ViTs where we control training conditions and isolate potential contributing factors. Specifically, we hypothesize that training and stochastic depth (huang2016deepnetworksstochasticdepth) may promote the emergence of block-recurrent patterns.
Motivated by evidence that residual networks tolerate variable effective depth (wu2019blockdropdynamicinferencepaths), we examined the effect of stochastic depth (SD) on block recurrence. During training, each layer is dropped independently with probability $p$, applied uniformly across depth. We trained ViT-B/14 from random initialization on CIFAR-100, using the cls token for the linear probe across a sweep of SD $p$ rates. We observe an increase in layer-layer similarity with increasing SD $p$ rates (Figure 4A, E). Next, we used these trained ViT networks as teachers for student Raptor  models (see Appendix A). Raptor  models were trained to reconstruct the hidden activation states of the ViT teacher across layers. Raptor forward passes are fully autoregressive, meaning each layer’s output is fed into the next layer and is also trained to match the corresponding layer in the teacher network (Figure 5). We quantify the similarity of the cls and patch token representations in each layer between the teacher and student networks as the $R^{2}$ of their matched token embeddings (Figure 4B).
@@ Results.
At scale, BRH holds in practice: a two-block Raptor recovers most of DINOv2 ViT-B, with three blocks essentially closing the gap. This reveals a strong simplicity bias in trained ViTs: depth reuses a small set of computations, effectively trading parameters for iterations. This reuse has two immediate consequences. First, it compresses the program description length that realizes the network’s computation, suggesting low algorithmic complexity. However, the implication is more subtle than standard Kolmogorov complexity (kolmogorov1965three). While Kolmogorov compression can replace a long program with an arbitrarily short one that runs in unbounded time, Raptor crucially preserves computational cost: applying the same block $n_{j}$ times achieves equivalent runtime to $n_{j}$ distinct untied blocks. In other words, ViTs admit a more compact program representation under the same runtime budget, aligning more closely with Levin’s complexity $K_{\text{Levin}}$ (levin1973universal).
@@
See Appendix E for details. Theoretically, BRH guarantee low Levin complexity (compact algorithmic descriptions at unchanged computational cost). Our empirical validation of BRH on DINOv2 then suggest that foundation vision models are algorithmically simpler than their nominal architecture suggests. This compression reinforces emerging evidence that simplicity principles govern successful neural networks (goldblum2023no; valle2018deep; huh2021low). For interpretability research, this offers an encouraging perspective: there exist representational lenses under which seemingly complex models reveal underlying simplicity. High-performing ViTs discover and iteratively reuse a compact set of algorithmic primitives, which is a structural regularity that may provide tractable entry points for mechanistic understanding.
@@ 4 From Block Recurrence to Dynamical Interpretability in VITs
Having observed block-structured representational similarity and confirmed that this similarity translates to functional recurrence, even in foundational models, we are naturally inclined to now seriously consider Vision Transformers as dynamical systems that can be interpreted using dynamical systems analysis tools schmid2022dynamic; ostrow2023beyond; huang2025inputdsa – what we term dynamical interpretability. We begin by establishing the basic dynamical properties of this depth flow, and present three key findings: (i) tokens converge directionally toward angular attractors with self-correcting dynamics, (ii) different token types exhibit specialized dynamics with punctuated transitions at phase boundaries, and (iii) later layers exhibit low-rank collective motion under weak contraction, reminiscent of mean-field processes with collapsing update dimensionality.

### v2

@@ Abstract
As Vision Transformers (ViTs) become standard backbones across vision, a mechanistic account of their computational phenomenology is now essential. Despite architectural cues that hint at dynamical structure, there is no settled framework that interprets Transformer depth as a well-characterized flow. In this work, we introduce the Block-Recurrent Hypothesis (BRH), arguing that trained ViTs admit a block-recurrent depth structure such that the computation of the original $L$ blocks can be accurately rewritten using only $k\ll L$ distinct blocks applied recurrently. Across diverse ViTs, between-layer representational similarity matrices suggest few contiguous phases. Yet, representational similarity does not necessarily imply functional similarity. To determine whether these phases reflect genuinely reusable computation, we operationalize our hypothesis in the form of block recurrent surrogates of pretrained ViTs, which we call Recurrent Approximations to Phase-structured TransfORmers (Raptor). Using small-scale ViTs, we demonstrate that phase-structure metrics correlate with our ability to accurately fit Raptor and identify the role of training and stochastic depth in promoting the recurrent block structure. We then provide an empirical existence proof for BRH in foundation models by showing that we can train a Raptor model to recover $96\%$ of DINOv2 ImageNet-1k linear probe accuracy in only 2 blocks while maintaining equivalent runtime. To provide a mechanistic account of these observations, we leverage our hypothesis to develop a program of Dynamical Interpretability. We find (i) directional convergence into class-dependent angular basins with self-correcting trajectories under small perturbations, (ii) token-specific dynamics, where cls executes sharp late reorientations while patch tokens exhibit strong late-stage coherence reminiscent of a mean-field effect and converge rapidly toward their mean direction, and (iii) a collapse of the update to low rank in late depth, consistent with convergence to low-dimensional attractors.
@@ 1 Introduction
In the last decade, Transformers have become the default neural network architecture across machine learning communities, scaling favorably with data and compute (Vaswani et al., 2017; Kaplan et al., 2020). In particular, Vision Transformers (ViTs) (Dosovitskiy et al., 2020) have become the core architecture used in visual foundation modeling frameworks such as DINOv2 (Oquab et al., 2023; Darcet et al., 2023) and CLIP (Radford et al., 2021); and have come to dominate a wide range of visual tasks, from general visual feature extraction (He et al., 2021; Chiu et al., 2024; Yun, 2025), to diffusion (Peebles and Xie, 2023), image segmentation (Kirillov et al., 2023; Liu et al., 2024), and video processing (Arnab et al., 2021; Baldassarre et al., 2025). This increasing breadth of use motivates a move from empirical optimization to principled understanding.
@@
Two pressures make this understanding urgent. First, safety-critical deployments (Wang and Chung, 2022; Alecu et al., 2022) demand mechanisms whose internal computation is inspectable (Losch et al., 2021), diagnosable (Adebayo et al., 2020), and verifiable (Tjeng and Tedrake, 2019) rather than opaque. As these models proliferate across domains, the ability to explain (Doshi-Velez and Kim, 2017; Gilpin et al., 2018; Kim et al., 2018), manipulate, and verify their behavior becomes increasingly essential. Second, from a scientific inference perspective (Cichy and Kaiser, 2019), understanding what makes these models work is essential for explaining their success. Their strong performance suggests they have discovered effective strategies, and identifying these strategies could advance our broader understanding of visual intelligence. Independently of any comparison to human vision, the goal is to uncover the principles that make these systems so effective.
@@
One promising path toward such understanding is to search for underlying simplicity. Multiple approaches explore different facets of this simplicity, whether in functional expressivity (Montúfar et al., 2014; Telgarsky, 2015; Serra et al., 2018; Balestriero and others, 2018), symmetry (Cohen and Welling, 2014; Olah et al., 2020), or computation (Wilson, 2025; Goldblum et al., 2023; Schmidhuber, 1997; Mingard et al., 2025). Discovering such simplicity principle should improve both development (Bronstein et al., 2017; Frankle and Carbin, 2019) and interpretability (Bereska and Gavves, 2024; Carvalho et al., 2019; Fel, 2024; Ghorbani et al., 2017; Fel et al., 2023; Smilkov et al., 2017; Sundararajan et al., 2017; Zeiler and Fergus, 2014; Templeton et al., 2024; Bricken et al., 2023). Depth offers a concrete place to look for this simplicity. Residual connections have long suggested a link to dynamical systems (Liao and Poggio, 2020; Veit et al., 2016; Greff et al., 2016; Boulch, 2017; Haber and Ruthotto, 2017), hinting at implicit recurrence even when layers have distinct parameters. This convergent evidence makes plausible a form of algorithmic parsimony (Ma et al., 2022) in which a small set of blocks may be reused across many layers, trading parameters for iterations. Related perspectives support this view (Dingle et al., 2020). More concretely, residual updates invite a discrete-time interpretation of depth (Chen et al., 2018; Chalvidal et al., 2020; Sander et al., 2022), attention induces coupled token dynamics (Lu et al., 2019; Geshkovski et al., 2023), and in language models contiguous block recurrence has been observed and exploited (Geiping et al., 2025; Fernando and Guitchounts, 2025; Dehghani et al., 2018; Tan et al., 2023). This renewed interest in recurrent computation (Jolicoeur-Martineau, 2025; Venkatraman et al., 2025) and algorithmic complexity (Shaw et al., 2025; Dingle et al., 2020) as frameworks for understanding neural network simplicity bias further motivates our investigation.
@@
However, no existing framework characterizes depth in ViTs as representational flow or determines whether apparent phases correspond to functional reuse. Furthermore, vision explainability research (Bach et al., 2015; Fong and Vedaldi, 2017; Novello et al., 2022; Muzellec et al., 2024; Petsiuk et al., 2018; Hedström et al., 2022; Fel et al., 2025; Gorton, 2024; Kowal et al., 2024; Bau et al., 2017; Vilas et al., 2023) has not leveraged dynamical systems analysis to model emergent network structure.
@@ 2 Emergent Phase Structure & the Block-Recurrent Hypothesis
Our investigation starts with a simple experiment: we construct layer-layer similarity matrices by computing the cosine similarity of each token at layer $l$ with the same token at layer $m$. As shown in Figure 1, despite significant differences in tasks, training objectives, and scale, all models exhibit consistent block-wise organization where contiguous layers exhibit high mutual similarity within blocks and lower similarity across block boundaries. This finding is not new and echoes early (and more recent) observations in residual networks (Kornblith et al., 2019; Hoang et al., 2025), but raises a fundamental question: does representational similarity reflect deeper computational structure? In fact, representational similarity alone provides no guarantee of functional equivalence. Layers might produce similar representations through entirely different computational pathways, or conversely, functionally equivalent computations might yield representations that appear dissimilar due to linear transformations or noise. The critical question is whether these apparent phases correspond to genuine functional recurrence – that is, whether the same computational operations are being reused across different layers within each phase. We formalize this possibility through the Block-Recurrent Hypothesis:
@@
However, recurrent architectures are notoriously difficult to train: rollouts can drift as small state errors compound over steps, and learning can be unstable because gradients propagated through many recurrent applications often vanish or explode (Pascanu et al., 2013; Linsley et al., 2020; De et al., 2024). Standard backpropagation through time becomes unstable as recurrence depth increases, and the model must learn to simultaneously handle both the forward dynamics and its own prediction errors in closed loop (Williams and Zipser, 1989). To circumvent these challenges, we will leverage the intermediate layer activations as training targets, enabling a staged approach that combines the stability of teacher forcing with the self-consistency required for autoregressive deployment. More precisely, we introduce a procedure to distill existing Vision Transformers into Recurrent Approximations to Phase-structured TransfORmers (Raptors), using $k$ parameter-tied blocks with repetition counts determined by a max-cut phase discovery algorithm. This approach transforms the abstract hypothesis into a concrete architectural and training framework that can be empirically validated.
This constructive approach requires that Raptor models reproduce the internal activations of the full ViT they approximate, similar to Dasgupta and Cohn (2025); Sanh et al. (2019); Shleifer and Rush (2020), not merely mimic the final output [^1] Unlike classical distillation, which typically supervises logits (and occasionally a few intermediate “hints”), we enforce one-to-one alignment of all layers representations across the entire depth for the same inputs. The recurrent surrogate must generate the teacher’s intermediate activations, not just its predictions.. The BRH implies that such reproduction should be possible within tolerance $\varepsilon$, making activation matching a natural training objective. Formally, let $\bm{f}$ be a reference ViT with intermediate activations $\bm{a}_{\ell}(\bm{x})\equiv\bm{f}_{\ell}(\bm{x})\in\mathbb{R}^{t\times d}$ for $\ell=0,\dots,L$, where layer $\ell=0$ denotes the patch encoder and $1\leq\ell\leq L$ refer to transformer layers. Here, $t$ is the number of tokens and $d$ the feature dimension. Let $\bm{B}_{j}$ denote the $j$-th parameter-tied block in our recurrent decomposition. The Raptor approximation produces activations:
@@
For a given number of blocks $k$, we must introduce a practical method to determine the number of recurrent iterations of each block ($n_{k}$); in other words, where the recurrent “phases” of computation begin and end. We accomplish this by casting this “block discovery” process as a weighted max-cut problem Goemans and Williamson (1995), solved via dynamic programming (see Subsection˜A.1 for details). Specifically, the algorithm seeks to partition depth into contiguous segments by maximizing within-block similarity and minimizing cross-block similarity. We visualize the results of this procedure applied to ViT-B in Figure˜3, demonstrating that the discovered blocks align reasonably with qualitative assessment.
@@
Having operationalized the BRH and demonstrated a method for block discovery, we now turn to the mechansitic origins of this phenomenon: under what conditions does this block-recurrent structure emerge in trained Vision Transformers? To investigate this systematically, we examine small-scale ViTs where we control training conditions and isolate potential contributing factors. Specifically, we hypothesize that training and stochastic depth (Huang et al., 2016) may promote the emergence of block-recurrent patterns.
Motivated by evidence that residual networks tolerate variable effective depth (Wu et al., 2019), we examined the effect of stochastic depth (SD) on block recurrence. During training, each layer is dropped independently with probability $p$, applied uniformly across depth. We trained ViT-B/14 from random initialization on CIFAR-100, using the cls token for the linear probe across a sweep of SD $p$ rates. We observe an increase in layer-layer similarity with increasing SD $p$ rates (Figure 4A, E). Next, we used these trained ViT networks as teachers for student Raptor  models (see Appendix A). Raptor  models were trained to reconstruct the hidden activation states of the ViT teacher across layers. Raptor forward passes are fully autoregressive, meaning each layer’s output is fed into the next layer and is also trained to match the corresponding layer in the teacher network (Figure 5). We quantify the similarity of the cls and patch token representations in each layer between the teacher and student networks as the $R^{2}$ of their matched token embeddings (Figure 4B).
@@ Results.
At scale, BRH holds in practice: a two-block Raptor recovers most of DINOv2 ViT-B, with three blocks essentially closing the gap. This reveals a strong simplicity bias in trained ViTs: depth reuses a small set of computations, effectively trading parameters for iterations. This reuse has two immediate consequences. First, it compresses the program description length that realizes the network’s computation, suggesting low algorithmic complexity. However, the implication is more subtle than standard Kolmogorov complexity (Kolmogorov, 1965). While Kolmogorov compression can replace a long program with an arbitrarily short one that runs in unbounded time, Raptor crucially preserves computational cost: applying the same block $n_{j}$ times achieves equivalent runtime to $n_{j}$ distinct untied blocks. In other words, ViTs admit a more compact program representation under the same runtime budget, aligning more closely with Levin’s complexity $K_{\text{Levin}}$ (Levin, 1973).
@@
See Appendix E for details. Theoretically, BRH guarantee low Levin complexity (compact algorithmic descriptions at unchanged computational cost). Our empirical validation of BRH on DINOv2 then suggest that foundation vision models are algorithmically simpler than their nominal architecture suggests. This compression reinforces emerging evidence that simplicity principles govern successful neural networks (Goldblum et al., 2023; Valle-Perez et al., 2019; Huh et al., 2023). For interpretability research, this offers an encouraging perspective: there exist representational lenses under which seemingly complex models reveal underlying simplicity. High-performing ViTs discover and iteratively reuse a compact set of algorithmic primitives, which is a structural regularity that may provide tractable entry points for mechanistic understanding.
@@ 4 From Block Recurrence to Dynamical Interpretability in VITs
Having observed block-structured representational similarity and confirmed that this similarity translates to functional recurrence, even in foundational models, we are naturally inclined to now seriously consider Vision Transformers as dynamical systems that can be interpreted using dynamical systems analysis tools Schmid (2022); Ostrow et al. (2023); Huang et al. (2025) – what we term dynamical interpretability. We begin by establishing the basic dynamical properties of this depth flow, and present three key findings: (i) tokens converge directionally toward angular attractors with self-correcting dynamics, (ii) different token types exhibit specialized dynamics with punctuated transitions at phase boundaries, and (iii) later layers exhibit low-rank collective motion under weak contraction, reminiscent of mean-field processes with collapsing update dimensionality.